# Multi-Agent Reinforcement Learning for Connected and Automated Vehicles Control: Recent Advancements and Future Prospects

### v1

@@ II-A1 Deep Q-Learning
Some widely recognized deep Q-learning algorithms include DQN [48], double deep Q-Network (DDQN)[28], and hindsight experience replay (HER) [49].
II-A2 Policy Gradient
In contrast to the Q-learning approach, the policy gradient method directly parameterizes the policy, represented as $\pi_{\theta}$, using a set of parameters $\theta$. The objective of updating $\theta$ is to enhance both the likelihood of taken actions and the cumulative rewards. This is accomplished through the loss function:
@@ II-A2 Policy Gradient
Subsequently, the policy network parameters undergo an update through stochastic gradient ascent, expressed as:
@@
Compared to Q-learning, policy gradient methods are robust to non-stationary transitions within each trajectory. However, it’s essential to note that they tend to exhibit high variance [32]. Renowned algorithms utilizing policy gradients include Trust Region Policy Optimization (TRPO) [45], Proximal Policy Optimization (PPO) [50], and Deep Deterministic Policy Gradient (DDPG) [43].
II-A3 Actor-critic Network
To address the high variance associated with sample returns in the policy gradient method, actor-critic algorithms, such as A2C [44], incorporate the advantage function to refine the policy gradient technique, leveraging both policy (actor) and value (critic) functions. The advantage function is represented as:
@@ II-A3 Actor-critic Network
Here, the parameters $\theta$ are updated through the policy loss function:
@@
Simultaneously, the value function updates with:
@@
where $\mathcal{D}$ represents the experience replay buffer, aggregating past experiences. This buffer works alongside parameters derived from previous iterations, typically used in a target network [35].
However, despite significant advancements, single-agent RL often struggles with scalability challenges, particularly in real-world control scenarios involving multiple agents. These challenges stem from inherent non-stationarities and the partial observability intrinsic to such systems [33].
II-B Multi-agent reinforcement learning (MARL)
In multi-agent systems involving multiple agents (e.g., AVs, see Fig. 3), the agent’s actions affect not only their own rewards but also the rewards of other agents. MARL has found great success across a wide range of multi-agent systems, including traffic light control[32], games [51], resource management in wireless networks [52], and power grid control [53], only to name a few. A prominent and simplistic methodology in this domain is Independent Q-learning (IQL) [54]. In IQL, each local Q-function predominantly relies on the local action, represented as $Q_{i}(s,a)\approx Q_{i}(s,a_{i})$. Along similar lines, the Independent Advantage Actor-Critic (IA2C) serves as an actor-critic version of MARL, as proposed by [4]. While both IQL and IA2C offer highly scalable solutions, they grapple with challenges posed by partial observability and non-stationary Markov Decision Processes (MDP). This is mainly attributed to their intrinsic assumption: the behaviors of all other agents are perceived as part of the environment’s dynamics. This becomes problematic considering these agents’ policies take continual updates during training [32].
@@ II-C1 Centralized training with decentralized execution (CTDE)
It is important to acknowledge the inherent trade-offs of the CTDE. One significant strength is the enhanced learning stability enabled by centralized training, which can result in more robust policies [59]. However, it is crucial to recognize that maintaining a centralized controller comes with its own set of challenges. Firstly, it can be prohibitively expensive and infeasible in certain scenarios, particularly in large-scale or resource-constrained environments. Additionally, the centralized controller introduces privacy concerns, as it necessitates the sharing of sensitive information among agents. Moreover, the central controller becomes a single point of failure, rendering the entire system vulnerable to disruptions. These trade-offs highlight the need for careful consideration of the CTDE framework’s suitability in specific MARL applications, taking into account the advantages and drawbacks it presents.
@@ II-C2 Decentralized training with decentralized execution (DTDE)
In decentralized settings, individual agents acquire knowledge independently, devoid of direct interactions. Each agent operates with its distinct set of observations, policies, and algorithms, utilizing the environment as the exclusive channel for engagement. Commonly employed algorithms within this framework encompass independent Q-learning (IQL) [54] and independent advantage actor-critic (IA2C) [4], frequently adopted as foundational benchmarks. Although these algorithms enable autonomous learning by agents, they may not fully harness the collaborative potential [60]. To address non-stationarity concerns, [61] proposes a fully decentralized MARL framework. In this framework, each agent makes autonomous decisions based on both its locally observed environment and the messages exchanged with neighboring agents via the network, offering convergence guarantees. Fig. 5 shows the DTDE pattern based on the actor-critic framework.
@@ II-D1 Value function decomposition
The problem of credit assignment in a fully cooperative environment has recently become a research hotspot. The shared rewards in a fully cooperative environment make it difficult to distinguish the contribution of each agent, some agents tend to be lazy [81]. To solve the above problem, some studies learn different value functions for each agent to distinguish the contribution of each agent. Foerster et al. [82] and Guo et al. [83] introduced the counterfactual baseline principle to learn different value functions by centralized learning method in the context of a shared team reward environment. Hou et al. utilized a role attention credit allocation mechanism to manage the role policy learning process in multi-agent systems [84]. Liu et al. construct the causal effect of the agent’s actions on the external state from the perspective of a causal diagram, solving the problem of lazy agents [78].
Among the above methods to solve credit assignment, value function decomposition is considered to be an important research direction. The main idea of value function decomposition is to decompose a joint value function into individual value functions to distinguish the different contributions of each agent. In the process, many useful value function methods were proposed to solve increasingly complex problems. VDN decomposed the joint action-value function into the sum of individual action-value functions for each agent, this linear decomposition method can achieve good results on some simple problems [62]. Different from the linear value function decomposition method of VDN, QMIX imposes monotonicity constraints between the joint value function and the individual value function, which enhances the representation ability of the joint value function and can deal with more and more complex scenarios [63]. QTRAN further improves the representation ability of joint functions by relaxing the non-negative weight network in QMIX [64]. WQMIX extends the value function method to complex scenarios of non-monotonic rewards by introducing weighting functions and non-monotonic mixing networks [65]. After that, some variant algorithms based on VDN and QMIX were used to solve different specific problems [66, 67].
@@ II-D2 Learning to Communicate
MARL research can be categorized into two groups based on their communication methods [33]. The first group operates without communication among agents and centers its efforts on stabilizing training by employing advanced value estimation methods. For example, MADDPG [85] extends DDPG to the MARL setting by employing a centralized critic network to enable the global value estimate. Similarly, COMA [86] extends the actor-critic algorithm to the MARL setting by estimating each advantage using a centralized critic and a counterfactual baseline. The second group explores heuristic communication protocols, which may involve either direct message sharing or learnable communication protocols. For instance, in FingerPrint [68], the low-dimensional policy fingerprints are directly shared and incorporated into the state space of neighboring agents. In DIAL [69], a communicated message is generated simultaneously with the estimation of action values by each DQN agent. This message is then encoded and aggregated with other input signals at the receiver side. NeurComm proposed in [33] argues that encoding and concatenating communicating signals, rather than aggregating them, offers distinct advantages in mitigating information loss during communication. IMAC method learns an efficient communication protocol as well as scheduling based on the information bottleneck principle [70]. Their method involves the development of a novel differentiable communication protocol, which incorporates information related to state, policy, and the encoded data from neighboring agents. Subsequently, this information is encoded and concatenated into the state space of the neighboring agents.
@@ II-D3 Hierarchical Structure
Feudal Multi-agent Hierarchies (FHM) establish a feudal hierarchy within a multi-agent context, yielding noteworthy results. However, a primary limitation of this approach is its inherent challenge in tackling fully collaborative tasks with shared rewards [71]. To address the predicament of multi-agent scenarios characterized by sparse and delayed rewards, several hierarchical structures with temporal abstraction have been proposed, including hierarchical QMIX and hierarchical communication network methods [72].
Furthermore, recognizing that certain hierarchical structures tend to heavily rely on manually crafted high-level action spaces, hierarchical learning with skill discovery (HSD) is designed to train a cooperative decentralized policy for high-level skill selection, coupled with the training of independent low-level policies for skill execution. This approach enhances adaptability to dynamic scenarios [73]. RODE, on the other hand, introduces a method for clear action space partitioning through action clustering, where each action corresponds to a distinct subspace, this innovation substantially mitigates the challenge of extensive parameter tuning associated in HSD [74]. Moreover, HAVEN introduces a dual coordination mechanism, encompassing inter-layer and inter-agent strategies, facilitated by the formulation of a reward function within a two-level hierarchical structure. This approach effectively addresses the instability issue through the concurrent optimization of strategies at all levels and inter-agent coordination [75]. Hierarchical multi-agent reinforcement learning is recognized as a promising avenue for addressing collaborative decision-making challenges in complex, large-scale scenarios. Nonetheless, it faces the challenge of difficult to design complex dynamic hierarchical structures and difficult migration of policies.
@@ II-D4 Causal Inference
There are multiple complex variables in multi-agent reinforcement learning, and it is usually difficult to directly discover their internal relationships. Therefore, some recent research has begun to combine MARL with causal inference to discover the causal relationships between agents or variables, and further understand the operating mechanism of agents through intervention and inference, thereby motivating the agent to conduct more targeted learning [91]. Pearl’s three-level causal model is considered a powerful tool for constructing causal relationships between variables [92]. Causal influence reward methods promote the collaborative performance of agents by rewarding those agents that have a causal influence on the actions of other agents, where this causal influence is evaluated by counterfactual reasoning [76]. Rafael Pina et al proposed the independent causal learning (ICL) algorithm to evaluate the causal effect of each agent’s observations on team collaboration performance and solve the credit allocation problem in the independent learning framework [77].
@@
[104], 2022 Bi-level DQN $\bullet$ An innovative reward function designed with their own benefits and the impact on overall traffic $\bullet$ The driving intentions of surrounding vehicles encoded into the observation space $\bullet$ Comfort $\bullet$ Safety $\bullet$ Travel efficiency
[105], 2022 MA2C $\bullet$ A novel local reward design and parameter-sharing $\bullet$ Multi-objective reward function design $\bullet$ Driving comfort $\bullet$ Safety $\bullet$ Travel efficiency
@@ III Towards applications of MARL in CAVs
CAVs have garnered substantial interest for their potential to reshape the transportation landscape, ushering in an era marked by heightened efficiency, enhanced safety, and bolstered sustainability [106]. Operating within networked environments, CAVs engage in interactions with multiple vehicles, promoting cooperative driving. For example, CAVs can acquire the skills to collaborate, execute seamless merges, and navigate intersections with precision, thereby advancing traffic safety and optimizing flow [106]. These transformative capabilities are increasingly being explored in tandem with MARL to further optimize CAV interactions within networked environments. Within the realm of MARL, CAVs are designed to engage in intelligent interactions with multiple vehicles, thus promoting cooperative driving strategies.
@@ III-A One-dimensional cooperation
As shown in Fig. 6, the platooning control problem is often formulated as a model-free multi-agent network [33], where each agent (i.e., AV) is capable of communicating with the vehicles ahead and behind via V2V communication channels, to achieve safe, fuel efficiency, and smooth vehicle following while maximizing the benefits of close formation driving. They learn cooperative strategies to adapt their behaviors based on the actions of other agents in the network.
@@
One critical task of platooning control is to achieve string stability, which refers to the ability of a line of CAVs to maintain a stable and orderly formation (e.g., stable distance and speed) as they travel in close proximity to each other [107]. Notably, MARL has been proposed to address and enhance the achievement of string stability in the platooning control scenario, contributing to the development of more efficient and coordinated platooning control strategies. For instance, in [95], a MARL approach based on a robust communication protocol with Long Short-Term Memory (LSTM) [108] is introduced, resulting in stable platooning control. In addition, Li et al. introduce the Communication Proximal Policy Optimization (CommPPO) algorithm to enhance stability, which adapts to varying agent numbers and supports different dynamics of platooning control. CommPPO incorporates a predecessor-leader-follower communication protocol to facilitate the transmission of both global and local state information among agents. Notably, CommPPO introduces a novel reward communication channel, effectively mitigating issues related to spurious rewards and mitigating the problem of “lazy agents”, which are commonly encountered in other MARL approaches [109]. In [33], the platooning control problem is formulated as a spatiotemporal Markov decision process. They enhance system stability by introducing a spatial discount factor for local agents. Furthermore, they introduce a novel differentiable communication protocol, NeurComm, which mitigates non-stationarity, leading to improved learning efficiency and control performance.
@@
Efficient communication among CAVs in platooning control also poses a fundamental challenge in achieving seamless coordination and maintaining the desired following properties, such as precise inter-vehicle spacing and synchronized maneuvers. For instance, Liu et al. introduce a Multi-Agent Hierarchical Attention Reinforcement Learning (MAHARL) framework, which incorporates the Graph Attention Network (GAT) at each hierarchical level to account for the mutual influences among agents. The hierarchical architecture of MAHARL empowers agents with foresight, enabling them to make favorable decisions, even in the absence of immediate incentives, and consider long-term rewards, as elaborated in their work [97]. Additionally, in scenarios where complete information is lacking, Li et al. [98] devise a game-theoretic approach to model the strategic interactions between the leading vehicle and the following vehicles. To overcome the absence of full information, the equilibrium solution in platooning control is identified through backward induction, coupled with information collection from the following vehicles.
@@
Mixed platooning control refers to the coexistence of multiple platoons. In this context, it is crucial for each platoon to possess comprehensive knowledge about the other existing platoons and vehicles in the network, including both inter-platooning connectivity, which pertains to communication between different platoons, and intra-platoon connectivity, which relates to communication between vehicles within a platoon [110]. Parvini et al. present two innovative MARL algorithms, namely, modified MADDPG and modified MADDPG with task decomposition. In these algorithms, each platoon leader acts as an independent agent, engaging with the environment to learn its optimal policy. Notably, these algorithms employ multiple critics to estimate both global and local rewards, thereby fostering cooperative behavior among the agents. In the second algorithm, individual rewards are further decomposed into task-specific sub-reward functions, offering insights into their work as detailed in [111].

### v3

@@ II-A1 Value-Based Methods
Some widely recognized deep Q-learning-based algorithms include DQN [47], double deep Q-Network (DDQN)[24], and hindsight experience replay (HER) [48].
II-A2 Policy-Based Methods
Diverging from the value-based Q-learning approach, policy-based methods prioritize direct manipulation of the policy itself, denoted as $\pi_{\theta}$, via a distinct set of parameters $\theta$. The primary goal of adjusting $\theta$ revolves around amplifying the likelihood of selected actions along the overall rewards accumulation. This process is guided by a specifically defined loss function, which is mathematically represented as:
@@ II-A2 Policy-Based Methods
Subsequently, the parameters of the policy network are updated by employing stochastic gradient ascent, facilitating an iterative improvement of the policy, as detailed by
@@
In contrast to Q-learning methods, policy gradient techniques demonstrate enhanced resilience to the non-stationary dynamics characterizing individual trajectories, albeit at the cost of potentially higher variance in outcomes [34]. Among the most prominent algorithms that embody the principles of policy gradients are Deep Deterministic Policy Gradient (DDPG) [40], Soft Actor-Critic (SAC) [49], and Asynchronous Advantage Actor-Critic (A3C) [41].
II-A3 Actor-Critic Methods
Actor-critic methods aim to mitigate the high variability in outcomes often seen with policy gradient techniques by integrating an advantage function. This function enhances the policy gradient methodology by utilizing both the policy update (actor) and a value estimation function (critic) [41]. The advantage function, a pivotal element in this approach, is defined as:
@@ II-A3 Actor-Critic Methods
where the update of parameters $\theta$ is guided by a policy loss function articulated as:
@@
Concurrently, the critic value function is refined through:
@@
with $\mathcal{D}$ denoting the experience replay buffer, a repository for history experiences. This buffer, in collaboration with parameters from preceding iterations often employed within a target network, facilitates the learning process [32].
Despite considerable advancements, single-agent RL algorithms frequently encounter scalability challenges, especially in complex, real-world physical environments that involve multiple agents. These difficulties stem from inherent non-stationarities and the partial observability characteristic on multi-agent scenarios [50].
II-B Preliminaries of MARL
In systems with multiple agents (e.g., AVs in Fig. 3), not only the reward agents themselves but also the reward of their neighbors will be influenced by their actions. MARL has been widely used in various complex systems, including managing traffic [34], strategic gameplay [51], optimizing wireless network resources [52], and configuring power grids [53], to highlight a few areas. A fundamental approach within this field is Independent Q-learning (IQL), where each agent’s Q-function is primarily developed based on local actions, as $Q_{i}(s,a)\approx Q_{i}(s,a_{i})$, simplifying the learning process. Similarly, the Independent Advantage Actor-Critic (IA2C) [45] represents a variant of MARL, indicating the breadth of strategies explored in this space However, both IQL and IA2C face challenges in applications due to the partial observability and dynamic environment, as they operate under the assumption that all other agents’ actions are part of the environment, which complicates learning when agents’ strategies continually evolve [34].
@@ II-B Preliminaries of MARL
The majority of applications reviewed in Section III of this paper mainly focus on cooperative scenarios, highlighting the importance of collaboration in achieving shared objectives in multi-agent systems.
@@ II-C1 Centralized training with decentralized execution (CTDE)
It is important to acknowledge the inherent trade-offs of the CTDE. One significant strength is the enhanced learning stability enabled by centralized training, which can result in more robust policies [58]. However, it is crucial to recognize that maintaining a centralized controller comes with its own set of challenges. Firstly, it can be prohibitively expensive and infeasible in certain scenarios, particularly in large-scale or resource-constrained environments. Additionally, the centralized controller introduces privacy concerns, as it necessitates the sharing of sensitive information among agents. Moreover, the central controller becomes a single point of failure, rendering the whole system vulnerable to disruptions. These trade-offs highlight the need for careful consideration of the CTDE framework’s suitability in specific MARL applications, taking into account the advantages and drawbacks it presents.
@@ II-C2 Decentralized training with decentralized execution (DTDE)
In decentralized settings, individual agents acquire knowledge independently, devoid of direct interactions. Each agent operates with its distinct set of observations, policies, and algorithms, utilizing the environment as the exclusive channel for engagement. Commonly employed algorithms within this framework encompass independent Q-learning (IQL) [59] and independent advantage actor-critic (IA2C) [45], frequently adopted as foundational benchmarks. Although these algorithms enable autonomous learning by agents, they may not fully harness the collaborative potential [60]. To address non-stationarity concerns, [61] proposes a fully decentralized MARL framework. In this framework, each agent makes decisions based on both its local observation and the messages exchanged with neighboring agents via the network, offering convergence guarantees. Fig. 5 shows the DTDE pattern based on the actor-critic framework.
@@ II-D1 Value function decomposition
The challenge of credit assignment in cooperative settings has emerged as a significant area of research interest. The shared rewards in a fully cooperative environment make it difficult to distinguish the contribution of each agent, some agents tend to be lazy [81]. To solve the above problem, some studies learn different value functions to distinguish the contribution of each agent. Foerster et al. [82] and Guo et al. [83] introduced the counterfactual baseline principle to learn different value functions by centralized learning method in the context of a shared team reward environment. Additionally, Hou et al. have developed a credit allocation mechanism based on role attention, which facilitates the learning process of role policies in multi-agent systems by managing how credit is allocated among agents [84]. Liu et al. construct the causal effect of the agent’s actions on the external state from the perspective of a causal diagram, solving the problem of lazy agents [78].
Among the above methods to solve credit assignment, value function decomposition is considered to be an important research direction. The main concept and feature of value function decomposition is to decompose a joint value function into individual components to identify the distinct contributions of each agent. In the process, many useful value function methods were proposed to solve increasingly complex problems. VDN decomposes the joint action-value function by breaking it down into a sum of the action-value functions for individual agents. This linear decomposition method yields favorable results in certain simple problems [62]. Different from the linear value function decomposition method of VDN, QMIX enhances the representational capacity of the joint value function by enforcing monotonicity constraints between joint and the individual value functions. This approach allows for handling increasingly complex scenarios effectively [63]. QTRAN further improves the representation ability of joint functions by relaxing the non-negative weight network in QMIX [64]. WQMIX extends the value function method to complex scenarios of non-monotonic rewards by introducing weighting functions and non-monotonic mixing networks [65]. After that, some variant algorithms based on VDN and QMIX were used to solve different specific problems [66, 67].
@@ II-D2 Learning to Communicate
Research in MARL can be categorized into two distinct groups according to their methods of communication [50]. The first group operates without communication among agents and centers its efforts on stabilizing training by employing advanced value estimation methods. For example, MADDPG [85] extends DDPG to the MARL setting by employing a centralized critic network to enable the global value estimate. Similarly, COMA [86] modifies the actor-critic method for use in MARL, calculating the advantage function for each agent by employing a centralized critic in conjunction with a counterfactual baseline. The second group investigates heuristic communication protocols, which may involve either direct message sharing or learnable communication protocols. For instance, in FingerPrint [68], the low-dimensional policy fingerprints are directly shared and incorporated into the state space of neighboring agents. In DIAL [69], each DQN agent concurrently produces a communicated message while assessing action values. This message is encoded and subsequently combined with additional input signals at the receiving end. NeurComm proposed in [50] argues that encoding and concatenating communicating signals, rather than aggregating them, offers distinct advantages in mitigating information loss during communication. IMAC method learns an efficient communication protocol as well as scheduling based on the information bottleneck principle [70]. Their method involves the development of a novel differentiable communication protocol, which incorporates information related to state, policy, and the encoded data from neighboring agents. Subsequently, this information is encoded and concatenated into the state space of the neighboring agents.
@@ II-D3 Hierarchical Structure
Feudal Multi-agent Hierarchies (FMH) establish a feudal hierarchy within a multi-agent context, yielding noteworthy results. However, a primary limitation of this approach is its inherent challenge in tackling fully collaborative tasks with shared rewards [71]. To address the predicament of multi-agent scenarios characterized by sparse and delayed rewards, several hierarchical structures with temporal abstraction have been proposed, including hierarchical QMIX and hierarchical communication network methods [72].
Furthermore, recognizing the limitations of traditional approaches, which typically rely on manually defined high-level action spaces, the Hierarchical Learning with Skill Discovery (HSD) model was conceived. This model aims to autonomously train decentralized policies for high-level skill selection while simultaneously developing independent low-level policies for the execution of these skills, thereby enhancing adaptability to changing scenarios [73]. On the other hand, learning roles to decompose (RODE) is introduced for clear action space partitioning through action clustering, where each action corresponds to a distinct subspace, this innovation substantially mitigates the challenge of extensive parameter tuning associated in HSD [74]. Moreover, HierArchical Value dEcompositioN (HAVEN) introduces a dual coordination mechanism, encompassing inter-layer and inter-agent strategies, facilitated by the formulation of a reward function within a two-level hierarchical structure. This approach effectively addresses the instability issue through the concurrent optimization of strategies at all levels and inter-agent coordination [75]. Hierarchical MARL is recognized as a promising avenue for addressing collaborative decision-making challenges in complex, large-scale scenarios. Nonetheless, it faces the challenge of difficulty in designing complex dynamic hierarchical structures and difficult migration of policies.
@@ II-D4 Causal Inference
There are multiple complex variables in MARL, and it is usually difficult to directly discover their internal relationships. Therefore, some recent research has begun to combine MARL with causal inference to discover the causal relationships between agents or variables, and further understand the operating mechanism of agents through intervention and inference, thereby motivating the agent to conduct more targeted learning [91]. Pearl’s three-level causal model is considered a powerful tool for constructing causal relationships between variables [92]. Causal influence reward methods promote the collaborative performance of agents by rewarding those agents that have a causal influence on the actions of other agents, where this causal influence is evaluated by counterfactual reasoning [76]. Rafael Pina et al proposed the independent causal learning (ICL) algorithm to evaluate the causal effect of each agent’s observations on team collaboration performance and solve the credit allocation problem in the independent learning framework [77].
@@
[104], 2022 Bi-level DQN $\bullet$ An innovative reward function designed with their own benefits and the impact on overall traffic $\bullet$ The driving intentions of the surrounding vehicles encoded into the observation space $\bullet$ Comfort $\bullet$ Safety $\bullet$ Travel efficiency
[105], 2022 MA2C $\bullet$ A innovative local reward and parameter-sharing $\bullet$ Multi-objective reward function $\bullet$ Driving comfort $\bullet$ Safety $\bullet$ Travel efficiency
@@ III Towards applications of MARL in CAVs
CAVs have garnered substantial interest for their potential to reshape the transportation landscape, ushering in an era marked by heightened efficiency, enhanced safety, and bolstered sustainability. Operating within networked environments, CAVs engage in interactions with multiple vehicles, promoting cooperative driving. For example, CAVs can acquire the capabilities to collaborate, execute seamless merges, and navigate intersections with precision, thereby advancing traffic safety and optimizing flow [106]. These transformative capabilities are increasingly being explored in tandem with MARL to further optimize CAV interactions within networked environments. Within the realm of MARL, CAVs are designed to engage in intelligent interactions with multiple vehicles, thus promoting cooperative driving strategies.
@@ III-A One-dimensional cooperation
As shown in Fig. 6, the problem of platooning control is commonly addressed through a model-free, multi-agent network approach [50]. In this framework, each agent, symbolizing an AV, has the ability to communicate with both preceding and following vehicles via vehicle-to-vehicle (V2V) communication channels. This setup aims to ensure safe, fuel-efficient, and smooth vehicle following operations, while also maximizing the advantages of driving in close formation [107].
Given $N$ vehicles in a platoon, the control objective is to maintain a desired spacing $L_{i}$ and ensure velocity matching. The state vector [108] for each vehicle $i$ at time $t$ is:
$\dot{\mathbf{x}}_{i}(t)=\left[\begin{array}[]{ll}0&1\\ 0&0\end{array}\right]\mathbf{x}_{i}(t)+\left[\begin{array}[]{l}0\\ 1\end{array}\right]\left(k_{p}\left(\left(x_{i-1}(t)-x_{i}(t)\right)-L_{i}\right)\right.$ (7)
$\left.+k_{d}\left(v_{i-1}(t)-v_{i}(t)\right)\right)$
where $x_{i}(t)$ is the position of the $ith$ vehicle at time $t$, $v_{i}(t)$ is the velocity of the $ith$ vehicle at time $t$, $L_{i}$ is the desired inter-vehicle distance (spacing) between vehicle $i$ and $i+1$, $\mathbf{x}_{i}(t)=\left[x_{i}(t),v_{i}(t)\right]^{T}$ is the state vector, $k_{p}$ and $k_{d}$ are the proportional gain for position error and the derivative gain for velocity error respectively. To optimize the platoon control, an objective function $J$ can be defined, incorporating both spacing error and velocity error over a time horizon $T$:
[…]
where $\alpha$ and $\beta$ are weighting factors that balance the importance of position and velocity errors. Thus, cooperative strategies would be developed to achieve the collective goal, adapting their behaviors in response to the actions of other vehicles within the network.
@@
One critical task of platooning control is to achieve string stability, which refers to the ability of a line of CAVs to maintain a stable and orderly formation (e.g., stable distance and speed) as they travel in close proximity to each other [109]. Notably, MARL has been proposed to address and enhance the achievement of string stability in the platooning control scenario, contributing to the development of more efficient and coordinated platooning control strategies. For instance, in [95], a MARL approach based on a robust communication protocol with Long Short-Term Memory (LSTM) [110] is introduced, resulting in stable platooning control. In addition, Li et al. introduce the Communication Proximal Policy Optimization (CommPPO) algorithm to enhance stability, which adapts to varying agent numbers and supports different dynamics of platooning control. CommPPO incorporates a predecessor-leader-follower communication protocol to facilitate the transmission of both global and local state information among agents. Notably, CommPPO introduces a novel reward communication channel, effectively mitigating issues related to spurious rewards and mitigating the problem of “lazy agents”, which are commonly encountered in other MARL approaches [111]. In [50], the platooning control problem is formulated as a spatiotemporal MDP. They enhance system stability by introducing a spatial discount factor for local agents. Furthermore, they introduce a novel differentiable communication protocol, NeurComm, which mitigates non-stationarity, leading to improved learning efficiency and control performance.
@@
Efficient communication among CAVs in platooning control also poses a fundamental challenge in achieving seamless coordination and maintaining the desired following properties, such as precise inter-vehicle spacing and synchronized maneuvers. For instance, Liu et al. introduce a Multi-Agent Hierarchical Attention RL (MAHARL) framework, which incorporates the Graph Attention Network (GAT) at each hierarchical level to account for the reciprocal influences between agents. The hierarchical architecture of MAHARL empowers agents with foresight, enabling them to make favorable decisions, even in the absence of immediate incentives, and consider long-term rewards, as elaborated in their work [97]. Additionally, in scenarios where complete information is lacking, Li et al. [98] develop a game-theoretic framework to capture the strategic interactions between the leading vehicle and the following vehicles. To overcome the absence of full information, the equilibrium solution in platooning control is identified through backward induction, coupled with information collection from the following vehicles.
@@
Mixed platooning control involves the simultaneous presence of multiple vehicle platoons within a network, necessitating a sophisticated level of coordination and communication both within (intra-platoon) and between (inter-platooning) these platoons. [112]. Parvini et al. have introduced two advanced MARL algorithms specifically designed to address the challenges of mixed platooning control. These include the modified MADDPG and its variant, the modified MADDPG with task decomposition. In these frameworks, the leaders of the platoons are treated as independent agents that interact with the environment to determine the most effective policies for platoon management. Notably, these algorithms employ multiple critics to estimate both global and local rewards, thereby fostering cooperative behavior among the agents. In the second algorithm, individual rewards are further decomposed into task-specific sub-reward functions, offering insights into their work as detailed in [113].